OpenStudy (anonymous):

The formula K equals C plus two hundred seventy-three shows how to estimate the temperature K measured in degrees Kelvin (°K) when you know the temperature C measured in degrees Celsius (°C). If room temperature is 295°K, what is room temperature in °C?

OpenStudy (cwrw238):

K = C + 273 when K = 295 we have 295 = C + 273 C is the room temperature in degrees Celsius can you find C from this equation?

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

nd no i can't imma little confused

OpenStudy (anonymous):

How do you undo addition?